DRAW 2 CONCENTRIC CIRCLES. tHE LARGER CIRCLE HAS A RADIUS THAT IS 18 INCHES LONGER THAN THE SMALLER CIRLCE. THE LARGER CIRCLE HAS A CIRCUMFERENCE THAT IS A PI FEET LARGER THAN THE SMALLER CIRCLE. FIND THE VALUE OF A.
Answer by mananth(16949)

smaller circle radius = r
larger circle radius = r+18
..
circumference of smaller circle = 2*pi*r
Circumference of larger circle = 2 * pi* (r+18)=
..
2pi(r+18)-2pir=A pi
2pi(r+18-r)=Api
divide by pi
2*18=A pi / pi
36 = A